How do I solve (2x-5)^2-180=0 Using the square root property?
Answer by stanbon(75887)

How do I solve (2x-5)^2-180=0 Using the square root property?
-----
(2x-5)^2 = 180
---
2x-5 = sqrt(180) or 2x-5 = -sqrt(180)
----
2x-5 = sqrt(36*5) or 2x-5 = -sqrt(36*5)
----
2x-5 = 6sqrt(5) or 2x - 5 = -6sqrt(5)
----
2x = 5+6sqrt(5) or 2x = 5-6sqrt(5)
-----
x = (5 + 6sqrt(5))/2 or x = (5-6sqrt(5))/2
==================
Cheers,
Stan H.
==================